A thin-plano convex lens acts like a concave mirror of 28cm focal length when is plane surface is silvered.Calculate the radius of curvature of its curved surface if the refractive index of the material of the lens s 1.5 and also find the power of the lens

Grade:10

1 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Ayush,

n1/u +n2/v=(n2-n1)/r

for first refraction by convex surface: u→∞, v=rn2/(n2-n1)

for plane reflection: r is infinite,v'=-v

for second refraction: 1.5/v' -1/v''=(1-1.5)/r

1/v''+1/v=1/f

power=1/f

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Ayush !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR